Hockeycheermom:

Find two rational expressions whose difference is equal - X+4/ 4x

dude:

\[\frac{4}{\left(x^2-4x\right)}-\frac{x}{\left(4x-16\right)}\]

Zepdrix:

SoccerMom, is that a negative in front of the x?

Zepdrix:

If you remember a basic fraction rule as I've noted above, then it might make this problem a bit easier.

Zepdrix:

Rearrange the terms in the numerator,\[\large\rm \frac{-x+4}{4x}\qquad=\qquad \frac{4+-x}{4x}\]and then apply the rule I mentioned in the picture.